In the reaction 2KClO3(s)= 2KCl(s)+ 3O2(g)

If the 50.0 L of oxygen at STP are needed, how many grams of KclO3 must be decomposed 


1
Expert's answer
2020-11-25T02:21:15-0500

Solution:

The balanced chemical equation:

2KClO3(s) = 2KCl(s)+ 3O2(g)

At STP, 1 mole of an ideal gas = 22.4 liters.

Thus:

50.0 L O2 × (1 mole/22.4 L) = 2.232 moles of O2


According to the chemical equation: n(KClO3)/2 = n(O2)/3

n(KClO3) = 2 × n(O2) / 3

n(KClO3) = (2 × 2.232 mol ) / 3 = 1.488 mol

Moles of KClO3 = 1.488 mol.


Moles of KClO3 = Mass of KClO3 / Molar mass of KClO3

The molar mass of KClO3 is 122.55 g mol-1.

Thus:

Mass of KClO3 = Moles of KClO3 × Molar mass of KClO3

Mass of KClO3 = 1.488 mol × 122.55 g mol-1 = 182.3544 g = 182.35 g

Mass of KClO3 is 182.35 grams = 182 grams (to 3 significant figures).


Answer: 182 grams of KClO3 must be decomposed.